OpenStudy (anonymous):

Solve: 3(x + 4) x - 2

OpenStudy (anonymous):

check your equaiton, you need an equal sign is it this? 3(x + 4) = x-2

OpenStudy (anonymous):

its an >= sign

OpenStudy (anonymous):

ok first you use the distributive property multipliy the 3 through all the terms in the parenthases

OpenStudy (anonymous):

\[3(x+4) \ge x-2\] \[3x + 12 \ge x -2\] now you subtract x and 12 from both sides \[2x \ge -14\] then you divide by 2 \[x \ge -7\] when you are dealing with inequalities you change the direction if you multiply or divide both sides of the equation by a negative number. We didn't have to do that in this problem

OpenStudy (anonymous):

thank you this was extremely helpful !

OpenStudy (anonymous):

no prob